2) What is the power consumed by the 2K 12 V = resistor? In {alhr AR? s 2K0 Tou Elov 31s 3.33 Kn
Answer #1

solution given that 4mA v A Yilma. 3 WAT Zaur . Ik - Ima lov using modal analysis ✓ + VLO -1=0 ++ y - 5 - 3r=12 Tv = 40. Takeak gesistog is Power consumed an PakravI FLIRI = 122 -( 3mA) 2 = 9x2 Pler.= 18mw.?